\providecommand{\href}[2]{#2}\begingroup\raggedright\begin{thebibliography}{10}

\bibitem{Jia:2024kbg}
L.-B. Jia, ``{A try for dark energy in quantum field theory: The vacuum energy of neutrino field},'' \href{http://arxiv.org/abs/2410.06604}{{\ttfamily arXiv:2410.06604 [hep-ph]}}.

\bibitem{Zheng:2024wkp}
S.-B. Zheng, ``{Neutrino oscillations originate from virtual excitation of Z bosons},'' \href{http://arxiv.org/abs/2407.00954}{{\ttfamily arXiv:2407.00954 [hep-ph]}}.

\bibitem{zheng2024neutrinos}
S.-B. Zheng, ``Neutrinos produced from $\beta$ decays of neutrons cannot be in coherent superpositions of different mass eigenstates,'' 2024.

\bibitem{Cline:2024iqp}
J.~M. Cline, ``{Comment on ''Neutrino oscillations originate from virtual excitation of $Z$ bosons'' and ''Neutrinos produced from $\beta$ decays of neutrons cannot be in coherent superpositions of different mass eigenstates''},'' \href{http://arxiv.org/abs/2410.05826}{{\ttfamily arXiv:2410.05826 [hep-ph]}}.

\bibitem{Addazi:2022kjt}
A.~Addazi, S.~Capozziello, Q.~Gan, and A.~Marcian\`o, ``{Dark energy and neutrino superfluids},'' \href{http://dx.doi.org/10.1016/j.dark.2022.101102}{{\em Phys. Dark Univ.} {\bfseries 37} (2022) 101102}, \href{http://arxiv.org/abs/2208.03591}{{\ttfamily arXiv:2208.03591 [hep-ph]}}.

\bibitem{Dey:2022cfg}
U.~K. Dey and U.~Sarkar, ``{Neutrino condensate dark energy from TeV scale extra dimensions},'' \href{http://dx.doi.org/10.1016/j.nuclphysb.2022.115768}{{\em Nucl. Phys. B} {\bfseries 979} (2022) 115768}, \href{http://arxiv.org/abs/2204.01141}{{\ttfamily arXiv:2204.01141 [hep-ph]}}.

\bibitem{Khalifeh:2020bdg}
A.~R. Khalifeh and R.~Jimenez, ``{Spinors and Scalars in curved spacetime: Neutrino dark energy (DE\ensuremath{\nu})},'' \href{http://dx.doi.org/10.1016/j.dark.2021.100777}{{\em Phys. Dark Univ.} {\bfseries 31} (2021) 100777}, \href{http://arxiv.org/abs/2010.08181}{{\ttfamily arXiv:2010.08181 [gr-qc]}}.

\bibitem{Gogoi:2020qif}
A.~Gogoi, R.~K. Sharma, P.~Chanda, and S.~Das, ``{Early Mass-varying Neutrino Dark Energy: Nugget Formation and Hubble Anomaly},'' \href{http://dx.doi.org/10.3847/1538-4357/abfe5b}{{\em Astrophys. J.} {\bfseries 915} no.~2, (2021) 132}, \href{http://arxiv.org/abs/2005.11889}{{\ttfamily arXiv:2005.11889 [astro-ph.CO]}}.

\bibitem{Salazar-Arias:2019xzx}
J.~G. Salazar-Arias and A.~P\'erez-Lorenzana, ``{Cosmological unification, dark energy, and the origin of neutrino mass},'' \href{http://dx.doi.org/10.1103/PhysRevD.101.083526}{{\em Phys. Rev. D} {\bfseries 101} no.~8, (2020) 083526}, \href{http://arxiv.org/abs/1907.00131}{{\ttfamily arXiv:1907.00131 [hep-ph]}}.

\bibitem{NobleChamings:2019ody}
F.~Noble~Chamings, A.~Avgoustidis, E.~J. Copeland, A.~M. Green, and B.~Li, ``{Early dark energy constraints on growing neutrino quintessence cosmologies},'' \href{http://dx.doi.org/10.1103/PhysRevD.100.043525}{{\em Phys. Rev. D} {\bfseries 100} no.~4, (2019) 043525}, \href{http://arxiv.org/abs/1904.00884}{{\ttfamily arXiv:1904.00884 [astro-ph.CO]}}.

\bibitem{Ennadifi:2017nvl}
S.~E. Ennadifi, ``{On neutrino masses and source of dark energy},'' \href{http://dx.doi.org/10.3103/S0027134917030055}{{\em Moscow Univ. Phys. Bull.} {\bfseries 72} no.~3, (2017) 271--273}.

\bibitem{Dey:2017wwt}
U.~K. Dey, T.~S. Ray, and U.~Sarkar, ``{Dark Energy from pNGB Mediated Dirac Neutrino Condensate},'' \href{http://dx.doi.org/10.1016/j.nuclphysb.2017.12.005}{{\em Nucl. Phys. B} {\bfseries 928} (2018) 258--267}, \href{http://arxiv.org/abs/1705.08484}{{\ttfamily arXiv:1705.08484 [hep-ph]}}.

\bibitem{McKellar:2014hia}
B.~H.~J. McKellar, T.~J. Goldman, and G.~J. Stephenson, ``{Effective masses in a dense fermion background - Applied to neutrinos, dark matter and dark energy},'' \href{http://dx.doi.org/10.1142/S0217751X14440102}{{\em Int. J. Mod. Phys. A} {\bfseries 29} (2014) 1444010}.

\bibitem{Guendelman:2012vc}
E.~I. Guendelman and A.~B. Kaganovich, ``{Neutrino generated dynamical dark energy with no dark energy field},'' \href{http://dx.doi.org/10.1103/PhysRevD.87.044021}{{\em Phys. Rev. D} {\bfseries 87} no.~4, (2013) 044021}, \href{http://arxiv.org/abs/1208.2132}{{\ttfamily arXiv:1208.2132 [gr-qc]}}.

\bibitem{Lambiase:2010ic}
G.~Lambiase, H.~Mishra, and S.~Mohanty, ``{Dark energy from Neutrinos and Standard Model Higgs potential},'' \href{http://dx.doi.org/10.1016/j.astropartphys.2012.02.003}{{\em Astropart. Phys.} {\bfseries 35} (2012) 629--633}, \href{http://arxiv.org/abs/1006.4461}{{\ttfamily arXiv:1006.4461 [hep-ph]}}.

\bibitem{Bhatt:2009wb}
J.~R. Bhatt, B.~R. Desai, E.~Ma, G.~Rajasekaran, and U.~Sarkar, ``{Neutrino Condensate as Origin of Dark Energy},'' \href{http://dx.doi.org/10.1016/j.physletb.2010.02.079}{{\em Phys. Lett. B} {\bfseries 687} (2010) 75--78}, \href{http://arxiv.org/abs/0911.5012}{{\ttfamily arXiv:0911.5012 [hep-ph]}}.

\bibitem{Ando:2009ts}
S.~Ando, M.~Kamionkowski, and I.~Mocioiu, ``{Neutrino Oscillations, Lorentz/CPT Violation, and Dark Energy},'' \href{http://dx.doi.org/10.1103/PhysRevD.80.123522}{{\em Phys. Rev. D} {\bfseries 80} (2009) 123522}, \href{http://arxiv.org/abs/0910.4391}{{\ttfamily arXiv:0910.4391 [hep-ph]}}.

\bibitem{Bhatt:2008dg}
J.~R. Bhatt, P.-H. Gu, U.~Sarkar, and S.~K. Singh, ``{Neutrino Dark Energy in Grand Unified Theories},'' \href{http://dx.doi.org/10.1103/PhysRevD.80.073013}{{\em Phys. Rev. D} {\bfseries 80} (2009) 073013}, \href{http://arxiv.org/abs/0812.1895}{{\ttfamily arXiv:0812.1895 [hep-ph]}}.

\bibitem{Antusch:2008hj}
S.~Antusch, S.~Das, and K.~Dutta, ``{Phenomenology of Hybrid Scenarios of Neutrino Dark Energy},'' \href{http://dx.doi.org/10.1088/1475-7516/2008/10/016}{{\em JCAP} {\bfseries 10} (2008) 016}, \href{http://arxiv.org/abs/0807.4930}{{\ttfamily arXiv:0807.4930 [astro-ph]}}.

\bibitem{Bamba:2008jq}
K.~Bamba, C.~Q. Geng, and S.~H. Ho, ``{Radiative neutrino mass generation and dark energy},'' \href{http://dx.doi.org/10.1088/1475-7516/2008/09/001}{{\em JCAP} {\bfseries 09} (2008) 001}, \href{http://arxiv.org/abs/0806.0952}{{\ttfamily arXiv:0806.0952 [hep-ph]}}.

\bibitem{Bjaelde:2008yd}
O.~E. Bjaelde and S.~Hannestad, ``{Neutrino Dark Energy With More Than One Neutrino Species},'' \href{http://dx.doi.org/10.1103/PhysRevD.81.063001}{{\em Phys. Rev. D} {\bfseries 81} (2010) 063001}, \href{http://arxiv.org/abs/0806.2146}{{\ttfamily arXiv:0806.2146 [astro-ph]}}.

\bibitem{Bhatt:2008hr}
J.~R. Bhatt and U.~Sarkar, ``{Majorana Neutrino Superfluidity and Stability of Neutrino Dark Energy},'' \href{http://dx.doi.org/10.1103/PhysRevD.80.045016}{{\em Phys. Rev. D} {\bfseries 80} (2009) 045016}, \href{http://arxiv.org/abs/0805.2482}{{\ttfamily arXiv:0805.2482 [hep-ph]}}.

\bibitem{Takahashi:2008epd}
R.~Takahashi, ``{Dark Energy and Supersymmetric Mass Varying Neutrinos},'' \href{http://dx.doi.org/10.24532/soken.116.1_A143}{{\em Soryushiron Kenkyu Electron.} {\bfseries 116} no.~1, (2008) A143--A145}.

\bibitem{Ichiki:2008rh}
K.~Ichiki and Y.-Y. Keum, ``{Neutrino Masses from Cosmological Probes in Interacting Neutrino Dark-Energy Models},'' \href{http://dx.doi.org/10.1088/1126-6708/2008/06/058}{{\em JHEP} {\bfseries 06} (2008) 058}, \href{http://arxiv.org/abs/0803.2274}{{\ttfamily arXiv:0803.2274 [hep-ph]}}.

\bibitem{Ichiki:2008st}
K.~Ichiki and Y.-Y. Keum, ``{Cosmological Signatures of the Interaction between Dark-Energy and Massive Neutrinos},'' \href{http://arxiv.org/abs/0803.3142}{{\ttfamily arXiv:0803.3142 [astro-ph]}}.

\bibitem{Bhatt:2007ah}
J.~R. Bhatt, P.-H. Gu, U.~Sarkar, and S.~K. Singh, ``{Left-right symmetric model of neutrino dark energy},'' \href{http://dx.doi.org/10.1016/j.physletb.2008.03.059}{{\em Phys. Lett. B} {\bfseries 663} (2008) 83--85}, \href{http://arxiv.org/abs/0711.2728}{{\ttfamily arXiv:0711.2728 [hep-ph]}}.

\bibitem{Sarkar:2007nt}
U.~Sarkar, ``{Leptogenesis, Dark Energy, Dark Matter and the neutrinos},'' \href{http://dx.doi.org/10.1063/1.2803796}{{\em AIP Conf. Proc.} {\bfseries 939} no.~1, (2007) 124--133}, \href{http://arxiv.org/abs/0712.2675}{{\ttfamily arXiv:0712.2675 [hep-ph]}}.

\bibitem{Bjaelde:2007ki}
O.~E. Bjaelde, A.~W. Brookfield, C.~van~de Bruck, S.~Hannestad, D.~F. Mota, L.~Schrempp, and D.~Tocchini-Valentini, ``{Neutrino Dark Energy -- Revisiting the Stability Issue},'' \href{http://dx.doi.org/10.1088/1475-7516/2008/01/026}{{\em JCAP} {\bfseries 01} (2008) 026}, \href{http://arxiv.org/abs/0705.2018}{{\ttfamily arXiv:0705.2018 [astro-ph]}}.

\bibitem{Keum:2007pq}
Y.-Y. Keum, ``{Probing for Dynamics of Dark-Energy in Mass Varying Neutrinos: Cosmic Microwave Background Radiation and Large Scale Structure},'' \href{http://dx.doi.org/10.1142/S0217732307025388}{{\em Mod. Phys. Lett. A} {\bfseries 22} (2007) 2131--2142}, \href{http://arxiv.org/abs/0705.2204}{{\ttfamily arXiv:0705.2204 [hep-ph]}}.

\bibitem{Gu:2007mi}
P.-H. Gu, H.-J. He, and U.~Sarkar, ``{Dirac neutrinos, dark energy and baryon asymmetry},'' \href{http://dx.doi.org/10.1088/1475-7516/2007/11/016}{{\em JCAP} {\bfseries 11} (2007) 016}, \href{http://arxiv.org/abs/0705.3736}{{\ttfamily arXiv:0705.3736 [hep-ph]}}.

\bibitem{Takahashi:2007xp}
R.~Takahashi and M.~Tanimoto, ``{Dark energy and neutrino model in SUSY: Remarks on active and sterile neutrinos mixing},'' \href{http://dx.doi.org/10.1142/S0218301307006861}{{\em Int. J. Mod. Phys. E} {\bfseries 16} (2007) 1529--1540}, \href{http://arxiv.org/abs/0704.0186}{{\ttfamily arXiv:0704.0186 [hep-ph]}}.

\bibitem{Gu:2007ps}
P.-H. Gu, H.-J. He, and U.~Sarkar, ``{Neutrino dark energy and baryon asymmetry from Higgs sector},'' \href{http://dx.doi.org/10.1016/j.physletb.2007.07.014}{{\em Phys. Lett. B} {\bfseries 653} (2007) 419--423}, \href{http://arxiv.org/abs/0704.2020}{{\ttfamily arXiv:0704.2020 [hep-ph]}}.

\bibitem{Sahu:2007uh}
N.~Sahu and U.~Sarkar, ``{Predictive model for dark matter, dark energy, neutrino masses and leptogenesis at the TeV scale},'' \href{http://dx.doi.org/10.1103/PhysRevD.76.045014}{{\em Phys. Rev. D} {\bfseries 76} (2007) 045014}, \href{http://arxiv.org/abs/hep-ph/0701062}{{\ttfamily arXiv:hep-ph/0701062}}.

\bibitem{Das:2006ht}
S.~Das and N.~Weiner, ``{Late Forming Dark Matter in Theories of Neutrino Dark Energy},'' \href{http://dx.doi.org/10.1103/PhysRevD.84.123511}{{\em Phys. Rev. D} {\bfseries 84} (2011) 123511}, \href{http://arxiv.org/abs/astro-ph/0611353}{{\ttfamily arXiv:astro-ph/0611353}}.

\bibitem{Hill:2006hj}
C.~T. Hill, I.~Mocioiu, E.~A. Paschos, and U.~Sarkar, ``{Neutrino Phenomenology, Dark Energy and Leptogenesis from pseudo-Nambu-Goldstone Bosons},'' \href{http://dx.doi.org/10.1016/j.physletb.2007.06.006}{{\em Phys. Lett. B} {\bfseries 651} (2007) 188--194}, \href{http://arxiv.org/abs/hep-ph/0611284}{{\ttfamily arXiv:hep-ph/0611284}}.

\bibitem{Takahashi:2006be}
R.~Takahashi and M.~Tanimoto, ``{Dark Energy and Neutrino Model in SUSY},'' \href{http://dx.doi.org/10.1063/1.2735269}{{\em AIP Conf. Proc.} {\bfseries 903} no.~1, (2007) 656--659}, \href{http://arxiv.org/abs/hep-ph/0610347}{{\ttfamily arXiv:hep-ph/0610347}}.

\bibitem{Hall:2006br}
L.~J. Hall, H.~Murayama, M.~Papucci, and G.~Perez, ``{Probing Dark Energy via Neutrino \textbackslash{}\& Supernova Observatories},'' \href{http://arxiv.org/abs/hep-ph/0607109}{{\ttfamily arXiv:hep-ph/0607109}}.

\bibitem{Ringwald:2006ks}
A.~Ringwald and L.~Schrempp, ``{Probing Neutrino Dark Energy with Extremely High-Energy Cosmic Neutrinos},'' \href{http://dx.doi.org/10.1088/1475-7516/2006/10/012}{{\em JCAP} {\bfseries 10} (2006) 012}, \href{http://arxiv.org/abs/astro-ph/0606316}{{\ttfamily arXiv:astro-ph/0606316}}.

\bibitem{Zanzi:2006xr}
A.~Zanzi, ``{Neutrino dark energy and moduli stabilization in a BPS braneworld scenario},'' \href{http://dx.doi.org/10.1103/PhysRevD.73.124010}{{\em Phys. Rev. D} {\bfseries 73} (2006) 124010}, \href{http://arxiv.org/abs/hep-ph/0603026}{{\ttfamily arXiv:hep-ph/0603026}}.

\bibitem{Guendelman:2006ji}
E.~I. Guendelman and A.~B. Kaganovich, ``{Exotic low density fermion states in the two measures field theory: Neutrino dark energy},'' \href{http://dx.doi.org/10.1142/S0217751X06032538}{{\em Int. J. Mod. Phys. A} {\bfseries 21} (2006) 4373--4406}, \href{http://arxiv.org/abs/gr-qc/0603070}{{\ttfamily arXiv:gr-qc/0603070}}.

\bibitem{Ma:2006mr}
E.~Ma and U.~Sarkar, ``{Connecting dark energy to neutrinos with an observable higgs triplet},'' \href{http://dx.doi.org/10.1016/j.physletb.2006.03.083}{{\em Phys. Lett. B} {\bfseries 638} (2006) 356--358}, \href{http://arxiv.org/abs/hep-ph/0602116}{{\ttfamily arXiv:hep-ph/0602116}}.

\bibitem{Gu:2005eq}
P.-H. Gu, X.-J. Bi, and X.-m. Zhang, ``{Dark energy and neutrino CPT violation},'' \href{http://dx.doi.org/10.1140/epjc/s10052-007-0217-7}{{\em Eur. Phys. J. C} {\bfseries 50} (2007) 655--659}, \href{http://arxiv.org/abs/hep-ph/0511027}{{\ttfamily arXiv:hep-ph/0511027}}.

\bibitem{Zhang:2005ywa}
X.-m. Zhang, ``{Dark energy and its interactions with neutrinos},'' \href{http://dx.doi.org/10.1063/1.2149667}{{\em AIP Conf. Proc.} {\bfseries 805} no.~1, (2005) 3--9}, \href{http://arxiv.org/abs/hep-ph/0510072}{{\ttfamily arXiv:hep-ph/0510072}}.

\bibitem{Li:2005zd}
H.~Li, B.~Feng, J.-Q. Xia, and X.~Zhang, ``{Supernova constraints on models of neutrino dark energy},'' \href{http://dx.doi.org/10.1103/PhysRevD.73.103503}{{\em Phys. Rev. D} {\bfseries 73} (2006) 103503}, \href{http://arxiv.org/abs/astro-ph/0509272}{{\ttfamily arXiv:astro-ph/0509272}}.

\bibitem{Fardon:2005wc}
R.~Fardon, A.~E. Nelson, and N.~Weiner, ``{Supersymmetric theories of neutrino dark energy},'' \href{http://dx.doi.org/10.1088/1126-6708/2006/03/042}{{\em JHEP} {\bfseries 03} (2006) 042}, \href{http://arxiv.org/abs/hep-ph/0507235}{{\ttfamily arXiv:hep-ph/0507235}}.

\bibitem{Afshordi:2005ym}
N.~Afshordi, M.~Zaldarriaga, and K.~Kohri, ``{On the stability of dark energy with mass-varying neutrinos},'' \href{http://dx.doi.org/10.1103/PhysRevD.72.065024}{{\em Phys. Rev. D} {\bfseries 72} (2005) 065024}, \href{http://arxiv.org/abs/astro-ph/0506663}{{\ttfamily arXiv:astro-ph/0506663}}.

\bibitem{Barbieri:2005gj}
R.~Barbieri, L.~J. Hall, S.~J. Oliver, and A.~Strumia, ``{Dark energy and right-handed neutrinos},'' \href{http://dx.doi.org/10.1016/j.physletb.2005.08.075}{{\em Phys. Lett. B} {\bfseries 625} (2005) 189--195}, \href{http://arxiv.org/abs/hep-ph/0505124}{{\ttfamily arXiv:hep-ph/0505124}}.

\bibitem{Brookfield:2005td}
A.~W. Brookfield, C.~van~de Bruck, D.~F. Mota, and D.~Tocchini-Valentini, ``{Cosmology with massive neutrinos coupled to dark energy},'' \href{http://dx.doi.org/10.1103/PhysRevLett.96.061301}{{\em Phys. Rev. Lett.} {\bfseries 96} (2006) 061301}, \href{http://arxiv.org/abs/astro-ph/0503349}{{\ttfamily arXiv:astro-ph/0503349}}.

\bibitem{Bi:2004ns}
X.-J. Bi, B.~Feng, H.~Li, and X.-m. Zhang, ``{Cosmological evolution of interacting dark energy models with mass varying neutrinos},'' \href{http://dx.doi.org/10.1103/PhysRevD.72.123523}{{\em Phys. Rev. D} {\bfseries 72} (2005) 123523}, \href{http://arxiv.org/abs/hep-ph/0412002}{{\ttfamily arXiv:hep-ph/0412002}}.

\bibitem{Peccei:2004sz}
R.~D. Peccei, ``{Neutrino models of dark energy},'' \href{http://dx.doi.org/10.1103/PhysRevD.71.023527}{{\em Phys. Rev. D} {\bfseries 71} (2005) 023527}, \href{http://arxiv.org/abs/hep-ph/0411137}{{\ttfamily arXiv:hep-ph/0411137}}.

\bibitem{Guendelman:2004hk}
E.~I. Guendelman and A.~B. Kaganovich, ``{Neutrino dark energy},'' \href{http://arxiv.org/abs/hep-th/0411188}{{\ttfamily arXiv:hep-th/0411188}}.

\bibitem{Kaplan:2004dq}
D.~B. Kaplan, A.~E. Nelson, and N.~Weiner, ``{Neutrino oscillations as a probe of dark energy},'' \href{http://dx.doi.org/10.1103/PhysRevLett.93.091801}{{\em Phys. Rev. Lett.} {\bfseries 93} (2004) 091801}, \href{http://arxiv.org/abs/hep-ph/0401099}{{\ttfamily arXiv:hep-ph/0401099}}.

\bibitem{Firouzjahi:2022xxb}
H.~Firouzjahi, ``{Cosmological constant problem on the horizon},'' \href{http://dx.doi.org/10.1103/PhysRevD.106.083510}{{\em Phys. Rev. D} {\bfseries 106} no.~8, (2022) 083510}, \href{http://arxiv.org/abs/2201.02016}{{\ttfamily arXiv:2201.02016 [gr-qc]}}.

\end{thebibliography}\endgroup

\documentclass[twocolumn,prd,nofootinbib,aps,prl,floats,floatfix,amsmath,amssymb,longbibliography,secnumarabic]{revtex4-1} %
\usepackage[english]{babel}

\usepackage[letterpaper,top=2cm,bottom=2cm,left=3cm,right=3cm,marginparwidth=1.75cm]{geometry}

\usepackage{amsmath}
\usepackage{graphicx}
\usepackage[colorlinks=true,citecolor=red]{hyperref}

\newcommand{\beq}{\begin{equation}}
\newcommand{\eeq}{\end{equation}}

\newcommand{\bea}{\begin{eqnarray}}
\newcommand{\eea}{\end{eqnarray}}

\begin{document}

\title{Comment on ``A try for dark energy in quantum field theory: The vacuum energy of neutrino field''}
\author{James M.\ Cline}
\email{jcline@physics.mcgill.ca}
\affiliation{McGill University Department of Physics \& Trottier Space Institute, 3600 Rue University, Montr\'eal, QC, H3A 2T8, Canada}
\affiliation{Niels Bohr International Academy,The Niels Bohr Institute,
Blegdamsvej 17, DK-2100 Copenhagen Ø, Denmark }

\begin{abstract}
It was recently suggested (\url{https://arxiv.org/pdf/2410.06604}) that the small value of the dark energy of the universe could be explained in terms of the scale of neutrino masses through a simple quantum field theoretic mechanism.  I clarify that the quantity computed is only a threshold correction to the vacuum energy at the low $m_\nu$ scale, and therefore cannot explain the magnitude of the vacuum energy.

\end{abstract}

\maketitle

Recently Ref.\ \cite{Jia:2024kbg} attempted to explain the small observed value of the dark energy density of the Universe in terms of the scale of neutrino masses, starting from the standard expression for their one-loop contribution 
summing the zero-point energies.  It frequently happens that surprising proposals, which sound implausible at the intuitive level, are difficult to immediately debunk because they are expressed in language that is not quite the mainstream parlance (see Refs.\ \cite{Zheng:2024wkp,zheng2024neutrinos,Cline:2024iqp} for an example).  Ref.\ \cite{Jia:2024kbg} is such a case.  The author starts with the quartically divergent expression for the zero-point energies from  a free fermion of mass $m$, and makes subtractions to obtain 
the Coleman-Weinberg formula that, as usual, depends on the renormalization scale,
\beq
    \delta V \sim -{m^4\over 64\pi^2} \ln(m^2/\mu^2)\,.
\eeq 
He  observes that if $m\sim 0.01$\,eV, like for the standard model neutrinos,
and if $\mu$ is slightly larger, then $\delta V$ has the right magnitude to match the observed vacuum energy density. (For previous papers that made a link between the neutrino mass and dark energy scales, see Refs.\ 
\cite{Addazi:2022kjt,
Dey:2022cfg,
Khalifeh:2020bdg,
Gogoi:2020qif,
Salazar-Arias:2019xzx,
NobleChamings:2019ody,
Ennadifi:2017nvl,
Dey:2017wwt,
McKellar:2014hia,
Guendelman:2012vc,
Lambiase:2010ic,
Bhatt:2009wb,
Ando:2009ts,
Bhatt:2008dg,
Antusch:2008hj,
Bamba:2008jq,
Bjaelde:2008yd,
Bhatt:2008hr,
Takahashi:2008epd,
Ichiki:2008rh,
Ichiki:2008st,
Bhatt:2007ah,
Sarkar:2007nt,
Bjaelde:2007ki,
Keum:2007pq,
Gu:2007mi,
Takahashi:2007xp,
Gu:2007ps,
Sahu:2007uh,
Das:2006ht,
Hill:2006hj,
Takahashi:2006be,
Hall:2006br,
Ringwald:2006ks,
Zanzi:2006xr,
Guendelman:2006ji,
Ma:2006mr,
Gu:2005eq,
Zhang:2005ywa,
Li:2005zd,
Fardon:2005wc,
Afshordi:2005ym,
Barbieri:2005gj,
Brookfield:2005td,
Bi:2004ns,
Peccei:2004sz,
Guendelman:2004hk,
Kaplan:2004dq,
Firouzjahi:2022xxb}.)  The author does not explain why one should choose this value of $\mu$, nor why we should ignore all the other particles of the standard model that could contribute similarly.  But it was a good try.

\bibliographystyle{utphys}
\bibliography{sample}

\end{document}